OpenStudy (anonymous):

Determine if the following equations are functions. Assume y is the dependent variable write "yes" or "no".

OpenStudy (anonymous):

Determine if the following equations are functions. Assume y is the dependent variable write "yes" or "no". y=3x^2__ 3(x^2+y^2)=1__ 3=4x^2+2xy-7___ (y-2)^2+4x=1___ 4x^3-2y=16___ y^3=x-1___ x=3y^2___ 2x+2y=7___

OpenStudy (anonymous):

you can tell if something is a function by plugging in x (input) values if two different input values have the same y (output) values, the equation is not a function